OpenStudy (anonymous):

Can I please have help? Urgent question Solve the differential equation by Undetermined Coefficients. y''+y=2xsinx I need a detailed walkthrough, cant find the particular solution.

OpenStudy (anonymous):

Okay, I'll assume you have the right homogeneous solution. As your guess, try \(y_p=A\sin x+B\cos x+Cx\sin x+Dx\cos x\).

OpenStudy (anonymous):

is that the correct form? I thought it was (Ax^2+Bx)sinx + (Cx^2+Dx)cosx

OpenStudy (anonymous):

Let's say, as an example, that you were given \(2t+3\). Would your guess be \(At^2+Bt\), or \(At+B\)?

OpenStudy (anonymous):

\[At+B\]

OpenStudy (anonymous):

Right. It's for the same reason that you would pick the guess solution I suggested. You could also try \((Ax^2+Bx+C)\sin x+(\cdots)\cos x\), but you'll likely find that the quadratic constants are 0.

OpenStudy (anonymous):

So what would Yp' and Yp'' look like. (using the correct form of Yp)

OpenStudy (anonymous):

\[Yp=A \cos x+Bsinx+Cxcosx+Dxsinx\] \[Yp'=-Asinx+Bcosx+Ccosx-Cxsinx+Dsinx-Dxsinx\] \[Yp''=-Acosx-Bsinx-Csinx-Csinx-Cxcosx+Dcosx-Dsinx-Dxcosx\]

OpenStudy (anonymous):

The last term in \(y_p'\) isn't right, and I think an error follows from that in \(y_p''\). \[\begin{align*} y_p&=A\sin x+B\cos x+Cx\sin x+Dx\cos x\\ y_p'&=A\cos x-B\sin x+C\sin x+Cx\cos x+D\cos x-Dx\sin x\\ &=(A+D)\cos x+(C-B)\sin x+Cx\cos x-Dx\sin x\\ y_p''&=-(A+D)\sin x+(C-B)\cos x+C\cos x-Cx \sin x-D\sin x-Dx\cos x\\ &=-(A+2D)\sin x+(2C-B)\cos x-Cx\sin x-Dx\cos x \end{align*}\]

OpenStudy (anonymous):

I dont get how Yp is \[\frac{ 1 }{ 2 }x^2\cos x+\frac{ 1 }{ 2 }x \sin x\]

OpenStudy (anonymous):

Hmm, looks like I might have to brush up on my guess solutions. Just to be safe, I'd use \[y_p=(Ax^2+Bx+C)\sin x+(Dx^2+Ex+F)\cos x\]